OpenStudy (anonymous):

Samantha knows that she needs $22,000 for a 10% down payment on a house she can afford. She found an investment that earns 2.75% interest compounding monthly. How much should she put in the account now, rounded to the nearest dollar, to ensure she has the down payment amount in 5 years?

OpenStudy (kropot72):

The amount A accrued on a Principal P after n payment periods at i interest per payment period is given by: \[\large A=P(1+i)^{n}\] Substituting the given values into this formula we get: \[\large 22000=P(1+\frac{0.0275}{12})^{60}\ ...........(1)\] Now you need to solve equation (1) for P to find the value of the required amount to be deposited.